OpenStudy (anonymous):

Determine whether the series ∞ E (-1)^n+1 (n+3)/(n +10) j=1 converges absolutely, converges conditionally, or diverges, showing all work

OpenStudy (anonymous):

For a series \[ \sum_1^\infty a_n \] to convege, then \[ \lim a_n=0 \] Do we violate this with the given series

OpenStudy (anonymous):

With the given series \[ a_n=(-1)^{n+1} \frac {n+3}{n+10}\\ |a_n|= \frac {n+3}{n+10}\\ \] What is the llimit \[ \lim_{n\to \infty} \frac {n+3}{n+10} \]

OpenStudy (anonymous):

The limit is 1, so \(a_n\) does not go to zero and the series diverges.

OpenStudy (anonymous):

do u get it

OpenStudy (anonymous):

Thanks, @eliassaab and thanks for checkin' @prowrestler